Find the vertex angle of an isosceles triangle with base angles of (4x-9) and 3x+2). Show your work.
Answer by ankor@dixie-net.com(12692)

Find the vertex angle of an isosceles triangle with base angles of (4x-9) and 3x+2).
:
In an isosceles triangle the base angles are equal, therefore:
4x - 9 = 3x + 2
4x - 3x = 2 + 9
x = 11
:
Find the angle:
4(11) - 9 = 35 degrees
:
Vertex = 180 - 2(35) = 110 degrees
:
:
Check solution:
180 - 2(3x+2) =
180 - 2(3(11) + 2) =
180 - 2(35) = 110